# Belief Dynamics Reveal the Dual Nature of In-Context Learning and Activation Steering

### v1

@@ Abstract
Large language models (LLMs) can be controlled at inference time through prompts (in-context learning) and internal activations (activation steering). Different accounts have been proposed to explain these methods, yet their common goal of controlling model behavior raises the question of whether these seemingly disparate methodologies can be seen as specific instances of a broader framework. Motivated by this, we develop a unifying, predictive account of LLM control from a Bayesian perspective. Specifically, we posit that both context- and activation-based interventions impact model behavior by altering its belief in latent concepts: steering operates by changing concept priors, while in-context learning leads to an accumulation of evidence. This results in a closed-form Bayesian model that is highly predictive of LLM behavior across context- and activation-based interventions in a set of domains inspired by prior work on many-shot in-context learning. This model helps us explain prior empirical phenomena—e.g., sigmoidal learning curves as in-context evidence accumulates—while predicting novel ones—e.g., additivity of both interventions in log-belief space, which results in distinct phases such that sudden and dramatic behavioral shifts can be induced by slightly changing intervention controls. Taken together, this work offers a unified account of prompt-based and activation-based control of LLM behavior, and a methodology for empirically predicting the effects of these interventions.
@@ 1 Introduction
Large Language Models (LLMs) have begun demonstrating increasingly impressive capabilities (Brown et al., 2020; Kaplan et al., 2020; Bubeck et al., 2023; Chang et al., 2024). However, reliable use of these systems in practical applications mandates the design of protocols that ensure generated outputs satisfy desirable properties—e.g., avoiding violent or harmful speech, sycophantic responses, or engagement with unsafe queries (Bai et al., 2022b; a; Anwar et al., 2024). To this end, prior work targeting inference-time control of model behavior has developed two broad methodologies: input-level interventions via In-Context Learning (ICL), where contexts such as questions, instructions, dialog, or sequences of input-output examples are used to condition model behavior (Brown et al., 2020; Liu et al., 2023; Wei et al., 2022; Bai et al., 2022b; a), and representation-level interventions via activation steering, where a model’s behavior is modulated by directly intervening on its hidden activations (Turner et al., 2024; Geiger et al., 2021; Templeton et al., 2024). Practical approaches to ICL often involve an informal process of prompt engineering through trial-and-error (White et al., 2023; Sahoo et al., 2024), whereas approaches to activation steering typically use ad-hoc datasets of contrasting pairs of examples (Turner et al., 2024; Marks and Tegmark, 2024).
To better understand the empirical success of these methods, recent theoretical work has begun exploring how input and representation-level interventions impact the distribution of generated outputs. Specifically, ICL has been framed as a form of Bayesian inference, where context modulates a space of hypotheses learned during pretraining (Xie et al., 2021; Bigelow et al., 2023; Wurgaft et al., 2025; Arora et al., 2024). Activation steering, on the other hand, has been argued to be a direct consequence of models learning to match the data distribution, which leads them to develop linear representations of concepts in particular layers (Park et al., 2024b; 2025b; Ravfogel et al., 2025; Arora et al., 2016). Given the shared goals of ICL and activation steering, it is plausible that there is a broader framework that helps formalize the notion of control in a probabilistic system, with seemingly disparate approaches, such as ICL and activation steering, acting as specific instances of this framework.
@@ 2.1 In-Context Learning
In-Context Learning (ICL), where an LLM learns from linguistic context, is often contrasted with in-weights learning, where an LLM learns during (pre)training by adjusting model weights (Chan et al., 2022; Reddy, 2023; Lampinen et al., 2024; Nguyen and Reddy, 2024). While ICL is traditionally framed as few-shot learning (Brown et al., 2020), wherein exemplars corresponding to a task are offered to a model in-context and the model is expected to perform the demonstrated task on a novel query, there is a broader spectrum of language model capabilities that fall under the category of in-context learning (Lampinen et al., 2024; Park et al., 2025a; 2024a), e.g., zero-shot learning of a novel language (Gemini Team, 2023; Bigelow et al., 2023; Akyürek et al., 2024) or optimization of a utility function (Von Oswald et al., 2023; Demircan et al., 2024; Yin et al., 2024).
ICL as Bayesian Inference
As argued by Xie et al. (2021); Bigelow et al. (2023); Panwar et al. (2024); Zhang et al. (2023); Min et al. (2022) and recently verified by Wurgaft et al. (2025); Park et al. (2024a); Raventós et al. (2024) in toy domains, different perspectives and phenomenology associated with ICL can be captured in a unifying, predictive framework by casting ICL as Bayesian inference. We build on this perspective by formalizing a Bayesian account of ICL in practical, large-scale settings. Specifically, following prior work, we define the distribution of model outputs $y$ conditioned on input context $x$ as inference over latent concepts $c$:
@@ 2.2 Activation Steering
Linear representation hypothesis and activation steering
It is unclear precisely why activation steering methods work. These methods are similar in nature to analogies in word vector algebra (Mikolov et al., 2013), as in the classic example king : queen :: man : woman, which can be represented in vector algebra as $v(\text{king})-v(\text{queen})=v(\text{man})-v(\text{woman})$. The Linear Representation Hypothesis (Park et al., 2024b; 2025b) formalizes this connection in terms of embedding representation $\lambda(x)$ and an unembedding representation $\gamma(y)$, where output behavior given an input p(y|x)p(y\ |\ x) is the softmax of the inner product: $p(y\ |\ x)\propto\exp\big(\ \lambda(x)^{\top}\gamma(y)\ \big)$. If each concept variable $Y(C=c)$ is defined as a set of elements, e.g., $\{\text{man},\text{king}\}\in c_{\text{male}}$ or $\{\text{woman},\text{queen}\}\in c_{\text{female}}$, concept vectors correspond to directions between an ordered pair of values $c$, e.g., $\text{male}\rightarrow\text{female}$ or $\text{English}\rightarrow\text{Russian}$. As Park et al. (2024b) show, if a model has learned to match the log posterior odds between a concept and its complement, that is, if $\frac{p(c|\lambda(x))}{p(c^{\prime}|\lambda(x))}=\frac{p(c)}{p(c^{\prime})}$, then all directions for increasing $p(c|x)$ will be parallel and correspond to the steering vector identified via methods like CAA. In what follows, we build on this argument to model the effects of activation steering.
@@ 3 Many-Shot In-Context Learning Experiments
Figure 2: Replication of many-shot ICL results in persona domains (Anil et al., 2024)
@@
Next, we provide further details about our many-shot ICL experiments. Experimental details not provided below appear in App. E. We used three harmful persona datasets previously used for many-shot jailbreaking (Anil et al., 2024; Arora et al., 2024), as well as two additional (non-harmful) persona datasets from the same collection (Perez et al., 2022)  [^1] https://github.com/anthropics/evals/tree/main/persona. The three harmful personas represent the “dark triad” of personality traits: Psychopathy, Machiavellianism, and Narcissism. Each of these represents a distinct set of properties that, if present in deployed LLMs, could present a risk of harm to users. The two additional personas we test, Subscribes to Moral Nihilism and Believes Life Has No Meaning, are categorized as types of “Moral Nihilism”. These personas are not necessarily harmful, but instead represent an arbitrary set of behaviors that are suppressed by post-training methods such as RLHF (Perez et al., 2022). As depicted in Fig. 1, the datasets consist of 1000 questions of the form Is the following statement something you would say? <statement> with two possible responses $y\in\{\text{Yes},\ \text{No}\}$, where half the statements have Yes as the persona-matching behavior $y^{(c)}$, and half have No as the persona-matching behavior. Context $x$ in these domains consists of a sequence of chat-formatted user/assistant exchanges. These persona datasets were chosen because of LLMs’ relatively fast learning rates with these datasets where we can observe the full sudden learning dynamics (Fig. 2), including both transition points and final plateau values, with fewer than 128 in-context examples (i.e. $|x|\leq 128$), and because behavior $p(y|x)$ can easily be measured by taking the LLM’s token logit probabilities for Yes and No.
@@
In the following section, we develop a theoretical framework for understanding how both ICL and activation steering operate in terms of updating beliefs in an LLM, and a belief dynamics model that implements this framework. Our framework makes three key predictions, and for each prediction we describe relevant empirical findings with Llama-3.1-8B and compare LLM behavior with that of our belief dynamics model (Eq 8). The analyses in our main text used Llama-3.1-8b-Instruct (Dubey et al., 2024), a capable model that can be accommodated with relatively modest compute requirements. We also tested two additional LLMs of similar scale, Qwen-2.5-7b-Instruct and Gemma-2-9b-Instruct (Appendix B). The results shown in Fig. 4, Fig. 6, and Fig. 7 represent held-out predictions using 10-fold cross-validation across magnitude values. Overall, we find a very high correlation between LLM probabilities and predictions on held-out data ($r=0.98$, averaged across our 5 domains).
Figure 3: Belief updating with concept vectors  (Left) From a representational perspective, we assume that the default behavior of an LLM (e.g. Neutral Persona $c^{\prime}$) and the target behavior (Target Persona $c$) correspond to concept vectors. In-context learning (blue) directs the initial belief state from $c^{\prime}$ to increasingly point towards $c$ as a function of the log number of shots $|x|$. Activation steering (orange) similarly directs the belief state towards $c$ as a function of steering magnitude. (Right) We offer a parallel Bayesian perspective that in-context learning ($x_{k}$) and activation steering ($v$) both operate by changing an LLM’s belief in latent concepts $c$. In our theory, in-context learning updates the posterior belief through the likelihood function $p(x|c)$ (where $p(c|x)\propto p(x|c)$) and activation steering intervenes on concept priors $p(c)\rightarrow p^{\prime}(c)$.
4 A Belief Dynamics Model of ICL and Steering
@@
$p(c|x)=\frac{p(c)\ p(x|c)}{p(c)\ p(x|c)+p(c^{\prime})\ p(x|c^{\prime})}=\frac{o(c|x)}{1+o(c|x)}=\sigma\left(\log o(c|x)\right).$ (3)
@@ 4.1 Context is evidence: dynamics of in-context learning
The likelihood term captures the relative evidence for $c$ vs. $c^{\prime}$ from $N$ in-context examples. To model the log-likelihood, we follow Goodman et al. (2008) by assuming a concept’s log-likelihood declines proportionally to the number of labels that do not correspond to the expected labels for the concept. Denoting $l_{i}$ as the label for in-context example $i$ and $y_{i}^{(c)}$ as the concept-consistent label, we write:
$\log p(x|c)\propto-|\{i\in\{1,\dots,N\}\ |\ l_{i}\neq y_{i}^{(c)}\}|.$
In our experiments, all labels will correspond to $c$, and hence $\log p(x|c)=0$ and $\log p(x|c^{\prime})\propto-N$. Thus, the likelihood function can be expected to accumulate evidence linearly with in-context examples $N$. However, previous work has observed that the log-probability of next-token predictions scales as a power law with context size (Anil et al., 2024; Liu et al., 2024a; Park et al., 2025a). To account for this scaling, we follow Wurgaft et al. (2025) and model evidence accumulation as sub-linear by multiplying the log-likelihood by a discount factor $\tau(N)$. Under power-law growth of likelihood, we can show $\tau(N)=N^{-\alpha}$ and hence log-Bayes factor scales with context-size as $\log\frac{p(x|c)}{p(x|c^{\prime})}\propto N^{1-\alpha}$ (see App. A). Then, assuming a direct mapping between the concept-consistent label $y^{(c)}$ and the concept $c$, the model’s probability for a concept-consistent answer is simply $p(c|x)$, yielding the following expression:
$p(y^{(c)}|x)=p(c|x)=\sigma\left(-\log o(c|x)\right)=\sigma\Big(-\log\frac{p(c)}{p(c^{\prime})}-\gamma N^{1-\alpha}\Big),$ (4)
@@
Prediction 1  Based on the functional form in Eq. 4, we should expect $p(y^{(c)}|x)$ to follow a sigmoidal trend as $N^{1-\alpha}$ accumulates.
@@ Results
Building on our replication of results by Anil et al. (2024), we now show a more precise form of the sudden learning trend: specifically, in Fig. 4, we predict and demonstrate that in-context learning dynamics follow a sigmoid curve as a function of $N^{1-\alpha}$. This trend is captured effectively by our belief dynamics model, which uses a likelihood function that scales sub-linearly. This sub-linearity helps explain the results of prior work as well, since plotting the posterior as a function of log number of in-context exemplars should also yield a sudden learning trend. Beyond offering the precise functional form of this trend, we also show that in-context learning dynamics change as a function of steering magnitude, where positive steering magnitudes lead to similar ICL dynamics with fewer in-context examples (i.e., shifting the ICL curve leftwards) and negative magnitudes have the opposite effect (shifting the curve rightwards).

### v2

@@ Abstract
Large language models (LLMs) can be controlled at inference time through prompts (in-context learning) and internal activations (activation steering). Different accounts have been proposed to explain these seemingly disparate methods, yet their shared goal of controlling model behavior raises the question of whether they are instances of a broader framework. We develop a unifying, predictive account of LLM control from a Bayesian perspective, positing that both context- and activation-based interventions impact model behavior by altering its belief in latent concepts: steering operates by changing concept priors, while in-context learning leads to an accumulation of evidence. This results in a closed-form Bayesian model that is highly predictive of LLM behavior across context- and activation-based interventions in a set of domains inspired by prior work on many-shot in-context learning. Our model explains prior empirical phenomena—e.g., sigmoidal learning curves as in-context evidence accumulates—while predicting novel ones—e.g., additivity of both interventions in log-belief space, which results in distinct phases such that sudden and dramatic behavioral shifts can be induced by slightly changing intervention controls. Taken together, this work offers a unified account of prompt-based and activation-based control of LLM behavior, and a methodology for empirically predicting the effects of these interventions.
Belief dynamics, In-context learning, Activation Steering, Interpretability
@@ 1 Introduction
Large Language Models (LLMs) have begun demonstrating increasingly impressive capabilities (Brown et al., 2020; Kaplan et al., 2020; Bubeck et al., 2023; Chang et al., 2024). However, reliable use of these systems in practical applications mandates the design of protocols that ensure generated outputs satisfy desirable properties—e.g., avoiding violent or harmful speech, sycophantic responses, or engagement with unsafe queries (Bai et al., 2022b, a; Anwar et al., 2024). To this end, prior work targeting inference-time control of model behavior has developed two broad methodologies: input-level interventions via In-Context Learning (ICL), where contexts such as questions, instructions, dialog, or sequences of input-output examples are used to condition model behavior (Brown et al., 2020; Liu et al., 2023; Wei et al., 2022; Bai et al., 2022b, a), and representation-level interventions via activation steering, where a model’s behavior is modulated by directly intervening on its hidden activations (Turner et al., 2024; Geiger et al., 2021; Templeton et al., 2024). Practical approaches to ICL often involve an informal process of prompt engineering through trial-and-error (White et al., 2023; Sahoo et al., 2024), whereas approaches to activation steering typically use ad-hoc datasets of contrasting pairs of examples (Turner et al., 2024; Marks and Tegmark, 2024).
To better understand the empirical success of these methods, recent theoretical work has begun exploring how input and representation-level interventions impact the distribution of generated outputs. Specifically, ICL has been framed as a form of Bayesian inference, where context modulates a space of hypotheses learned during pretraining (Xie et al., 2021; Bigelow et al., 2023; Wurgaft et al., 2025; Arora et al., 2024). Activation steering, on the other hand, has been argued to be a direct consequence of models learning to match the data distribution, which leads them to develop linear representations of concepts in particular layers (Park et al., 2024b, 2025b; Ravfogel et al., 2025; Arora et al., 2016). Given the shared goals of ICL and activation steering, it is plausible that there is a broader framework that helps formalize the notion of control in a probabilistic system, with seemingly disparate approaches, such as ICL and activation steering, acting as specific instances of this framework.
@@ 2.1 In-Context Learning
In-Context Learning (ICL), where an LLM learns from linguistic context, is often contrasted with in-weights learning, where an LLM learns during (pre)training by adjusting model weights (Chan et al., 2022; Reddy, 2023; Lampinen et al., 2024; Nguyen and Reddy, 2024). While ICL is traditionally framed as few-shot learning (Brown et al., 2020), wherein exemplars corresponding to a task are offered to a model in-context and the model is expected to perform the demonstrated task on a novel query, there is a broader spectrum of language model capabilities that fall under the category of in-context learning (Lampinen et al., 2024; Park et al., 2025a, 2024a), e.g., zero-shot learning of a novel language (Gemini Team, 2023; Bigelow et al., 2023; Akyürek et al., 2024) or optimization of a utility function (Von Oswald et al., 2023; Demircan et al., 2024; Yin et al., 2024).
As argued by Xie et al. (2021); Bigelow et al. (2023); Panwar et al. (2024); Zhang et al. (2023); Min et al. (2022) and recently verified by Wurgaft et al. (2025); Park et al. (2024a); Raventós et al. (2024) in toy domains, different perspectives and phenomenology associated with ICL can be captured in a unifying, predictive framework by casting ICL as Bayesian inference. We build on this perspective by formalizing a Bayesian account of ICL in practical, large-scale settings. Following prior work, we define the distribution of model outputs $y$ conditioned on input context $x$ as inference over latent concepts $c$:
@@ 2.2 Activation Steering
Linear representation hypothesis
It is unclear precisely why activation steering methods work. These methods are similar in nature to analogies in word vector algebra (Mikolov et al., 2013), as in the classic example king : queen :: man : woman, which can be represented in vector algebra as $v(\text{king})-v(\text{queen})=v(\text{man})-v(\text{woman})$. The Linear Representation Hypothesis (Park et al., 2024b, 2025b) formalizes this connection in terms of embedding representation $\lambda(x)$ and an unembedding representation $\gamma(y)$, where output behavior given an input p(y|x)p(y\ |\ x) is the softmax of the inner product: $p(y\ |\ x)\propto\exp\big(\ \lambda(x)^{\top}\gamma(y)\ \big)$. If each concept variable $Y(C=c)$ is defined as a set of elements, e.g., $\{\text{man},\text{king}\}\in c_{\text{male}}$ or $\{\text{woman},\text{queen}\}\in c_{\text{female}}$, concept vectors correspond to directions between an ordered pair of values $c$, e.g., $\text{male}\rightarrow\text{female}$ or $\text{English}\rightarrow\text{Russian}$. As Park et al. (2024b) show, if a model has learned to match the log posterior odds between a concept and its complement, that is, if $\frac{p(c|\lambda(x))}{p(c^{\prime}|\lambda(x))}=\frac{p(c)}{p(c^{\prime})}$, then all directions for increasing $p(c|x)$ will be parallel and correspond to the steering vector identified via methods like CAA. In what follows, we build on this argument to model the effects of activation steering.
@@ 3 Many-Shot In-Context Learning Experiments
Figure 2: Replication of many-shot ICL results of Anil et al. (2024)
Figure 3: Belief updating with concept vectors    (Left) From a representational perspective, we assume that the default behavior of an LLM (e.g. Neutral Persona $c^{\prime}$) and the target behavior (Target Persona $c$) correspond to concept vectors. In-context learning (blue) directs the initial belief state from $c^{\prime}$ to increasingly point towards $c$ as a function of the log number of shots $|x|$. Activation steering (orange) similarly directs the belief state towards $c$ as a function of steering magnitude. (Middle, Right) We offer a parallel Bayesian perspective that in-context learning ($x_{k}$) and activation steering ($v$) both operate by changing an LLM’s belief in latent concepts $c$. In our theory, in-context learning updates the posterior belief through the likelihood function $p(x|c)$ (where $p(c|x)\propto p(x|c)$) and activation steering intervenes on concept priors $p(c)\rightarrow p^{\prime}(c)$.
@@
Next, we provide further details about our many-shot ICL experiments. Experimental details not provided below appear in App. H as well as Apps. EF. For our main experiments, we used three harmful persona datasets previously used for many-shot jailbreaking (Anil et al., 2024; Arora et al., 2024), as well as two additional (non-harmful) persona datasets from the same collection (Perez et al., 2022). The three harmful personas represent the “dark triad” of personality traits: Psychopathy, Machiavellianism, and Narcissism. Each of these represents a distinct set of properties that, if present in deployed LLMs, could present a risk of harm to users. The two additional personas we test, Subscribes to Moral Nihilism and Believes Life Has No Meaning, are categorized as types of “Moral Nihilism”. These personas are not necessarily harmful, but instead represent an arbitrary set of behaviors that are suppressed by post-training methods such as RLHF (Perez et al., 2022). We also analyzed four additional concepts which represent an aribtrary set of concepts with relatively higher priors in LLMs: Machiavellianism, Interest in Science, Interest in Music, Desire to Create Allies (App. E). As depicted in Fig. 1, these datasets consist of 1000 questions of the form Is the following statement something you would say? <statement> with two possible responses $y\in\{\text{Yes},\ \text{No}\}$, where half the statements have Yes as the persona-matching behavior $y^{(c)}$, and half have No as the persona-matching behavior. Context $x$ in these domains consists of a sequence of chat-formatted user/assistant exchanges. These persona datasets were chosen because of LLMs’ relatively fast learning rates with these datasets where we can observe the full sudden learning dynamics (Fig. 2), including both transition points and final plateau values, with fewer than 128 in-context examples (i.e. $|x|\leq 128$), and because behavior $p(y|x)$ can easily be measured by taking the LLM’s token logit probabilities for Yes and No. To demonstrate the generality of our approach, we also evaluated one dataset that did not use personas: flipped-label sentiment analysis, adopted from Agarwal et al. (2024). For each example datum in this task, the LLM is given a single sentence from the financial phrasebank dataset (Malo et al., 2014b) and is tasked with assigning one of three possible labels $y\in\{\text{Positive},\text{Neutral},\text{Negative}\}$, where labels permuted from their original meanings so that the LLM must learn a new mapping between sentiment labels and their underlying meanings (details in App. F).
@@
In the following section, we develop a theoretical framework for understanding how both ICL and activation steering operate in terms of updating beliefs in an LLM, and a belief dynamics model that implements this framework. Our framework makes three key predictions, and for each prediction we describe relevant empirical findings with Llama-3.1-8B and compare LLM behavior with that of our belief dynamics model (Eq 9). The analyses in our main text used Llama-3.1-8b-Instruct (Dubey et al., 2024), a capable model that can be accommodated with relatively modest compute requirements. We also tested two additional LLMs of similar scale, Qwen-2.5-7b-Instruct and Gemma-2-9b-Instruct (Appendix B). The results shown in Fig. 4, Fig. 6, Fig. 7, and Fig. 8 represent held-out predictions using 10-fold cross-validation across magnitude values. Overall, we find a very high correlation between LLM probabilities and predictions on held-out data ($r=0.98$, averaged across our 5 domains; $p<.001$ for all correlations).
4 Belief Dynamics Model of ICL and Steering
@@
$p(c|x)$ $=\frac{p(c)\ p(x|c)}{p(c)\ p(x|c)+p(c^{\prime})\ p(x|c^{\prime})}$
$=\frac{o(c|x)}{1+o(c|x)}=\sigma\left(\log o(c|x)\right).$ (3)
@@ 4.1 Context is evidence: dynamics of in-context learning
The likelihood term captures the relative evidence for $c$ vs. $c^{\prime}$ from $N$ in-context examples. To model the log-likelihood, we follow Goodman et al. (2008) by assuming a concept’s log-likelihood declines proportionally to the number of labels that do not correspond to the expected labels for the concept. Denoting $l_{i}$ as the label for in-context example $i$ and $y_{i}^{(c)}$ as the concept-consistent label (i.e. the behavior $y_{i}$ that is consistent with a concept $c$ rather than $c^{\prime}$), we write:
$\log p(x|c)\propto-|\{i\in\{1,\dots,N\}\ |\ l_{i}\neq y_{i}^{(c)}\}|.$ (4)
In our experiments, all labels will correspond to $c$, and hence $\log p(x|c)=0$ and $\log p(x|c^{\prime})\propto-N$. Thus, the likelihood function can be expected to accumulate evidence linearly with in-context examples $N$. However, previous work has observed that the log-probability of next-token predictions scales as a power law with context size (Anil et al., 2024; Liu et al., 2024b; Park et al., 2025a). To account for this scaling, we follow Wurgaft et al. (2025) and model evidence accumulation as sub-linear by multiplying the log-likelihood by a discount factor $\tau(N)$. Under power-law growth of likelihood, we can show $\tau(N)=N^{-\alpha}$ and hence log-Bayes factor scales with context-size as $\log\frac{p(x|c)}{p(x|c^{\prime})}\propto N^{1-\alpha}$ (see App. A). Then, assuming a direct mapping between the concept-consistent label $y^{(c)}$ and the concept $c$, the model’s probability for a concept-consistent answer is simply $p(c|x)$, yielding the following expression:
$p(y^{(c)}|x)$ $=p(c|x)=\sigma\left(-\log o(c|x)\right)$
$=\sigma\Big(-\log\frac{p(c)}{p(c^{\prime})}-\gamma N^{1-\alpha}\Big),$ (5)
@@
Prediction 1  Based on the functional form in Eq. 4.1, we should expect $p(y^{(c)}|x)$ to follow a sigmoidal trend as $N^{1-\alpha}$ accumulates.
@@ Results
Building on our replication of results by Anil et al. (2024), we now show a more precise form of the sudden learning trend: specifically, in Fig. 4, we predict and demonstrate that in-context learning dynamics follow a sigmoid curve as a function of $N^{1-\alpha}$. This trend is captured effectively by our belief dynamics model, which uses a likelihood function that scales sub-linearly. Note that in some cases, such as high-prior concepts (App. E), belief increases more sharply or plateaus earlier as a function of $N^{1-\alpha}$. This sub-linearity helps explain the results of prior work as well, since plotting the posterior as a function of log number of in-context exemplars should also yield a sudden learning trend. Beyond offering the precise functional form of this trend, we also show that in-context learning dynamics change as a function of steering magnitude, where positive steering magnitudes lead to similar ICL dynamics with fewer in-context examples (i.e., shifting the ICL curve leftwards) and negative magnitudes have the opposite effect (shifting the curve rightwards).